### v2

@@ 2 BACKGROUND
This section is intended to provide a brief background about ORAN, NTNs, and AI-Native for communication. We have also listed some important and representative survey or magazine papers in each subsection for the interested readers to further explore these topics.
2.1 ORAN
In this section, we aim to provide the readers with a clear understanding of the ORAN architecture and its key characteristics. A wide range of survey and tutorial articles has examined ORAN from various perspectives, providing valuable insights into its architecture, design principles, and implementation. Representative works [openRAN-4, openRAN-TN-1, azariah2024survey, openRAN-7, openRAN-8] offered comprehensive overviews of ORAN development, interfaces, and open research challenges. In addition, other surveys and tutorials presented a broad overview of ORAN along with focused discussions on specific topics such as network slicing [ORAN-survey-2025], simulation and testbeds [herrera2025tutorial], and spectrum allocation [martian2025towards]. The aforementioned surveys and tutorials can serve as valuable complementary resources for readers seeking further background on ORAN relevant to this study.
2.1.1 ARCHITECTURE
The RAN architectures have evolved from the conventional ones in 2-4 G networks to virtualized RAN (vRAN) and cloud-RAN (C-RAN), where the network functions and services are decoupled from the hardware and carried out virtually [openRAN-3]. This evolution has further led to ORAN, where multi-vendor components are disaggregated, virtualized, and connected through open and well-defined interfaces.
The O-RAN Alliance was founded in 2018 to promote openness and intelligence in the RAN after the two organizations C-RAN Alliance and the xRAN forum joined forces. The main goal of the Alliance is to standardize an architecture and set of interfaces for the ORAN paradigm [openRAN-2]. The alliance introduces the ORAN architecture and specifications to extend the RAN standards to include openness and intelligence. The alliance has 11 technical work groups (WGs) for the specification tasks in the ORAN architecture, 5 focus groups for the topics that are over-arching the technical WGs or are relevant for the whole organization, and a next Generation Research Group (nGRG) focusing on the research of open and intelligent RAN principles in 6G and future network standards [openRAN-1].
@@ 2.1.1 ARCHITECTURE
  • SMO: Service management and orchestration (SMO) provides the end-to-end lifecycle management, such as the orchestration and fault, configuration, accounting, performance, and security (FCAPS) operations, of the ORAN network functions and RICs [baranda2025architectural].
  • O-eNB: O-eNB may take the form of either a legacy eNB or an ng-eNB, since the ORAN architecture also supports the integration of LTE base stations.
@@
  • Y1 Consumer: Y1 Consumer refers to the entity within or outside of the public land mobile network (PLMN) trust domain that consumes the Y1 services produced by the near-real-time RIC (Near-RT RIC) [oran_y1_interface_2024].
ORAN architecture defines disaggregated functional entities, including the ORAN radio unit (O-RU), ORAN distributed unit (O-DU), and ORAN centralized unit (O-CU), which are interconnected through standardized open interfaces specified by the O-RAN Alliance and 3GPP. The O-RAN Alliance and 3GPP has specified different split options as discussed later in Section 4. Following the typical functional split Option 7.2x [larsen2018survey, oran_architecture_2025], the main functionalities of these entities are:
@@
In addition, two types of RICs, including non-real-time RIC (Non-RT RIC) and Near-RT RIC, enable the intelligent RAN optimization across different layers and timescales [oranwhitepaper-1, oran_architecture_2025]. Their key functionalities are summarized as follows:
@@
To provide essential background on the interfaces defined by 3GPP but adopted by the O-RAN Alliance [oran_architecture_2025, ORAN-survey-2025], brief descriptions are presented as follows:
@@
Figure 2: The latest ORAN Architecture by O-RAN Alliance [oran_architecture_2025].
Furthermore, we introduce the new interfaces defined by the O-RAN Alliance below [openRAN-4, oran_architecture_2025].
@@
  • Y1: Y1 interface enables the Y1 consumers to subscribe or request the RAN analytics information provided by Near-RT RIC [oran_y1_interface_2024].
  • D2: D2 interface is newly defined as the interface between two O-DU connected to the same O-CU-CP, and supports NR carrier aggregation of carriers configured in different O-DU nodes [oran_D2_interface_2025].
2.1.2 KEY FEATURES
@@
  • Open Interfaces: The open interface design enables full interoperability and flexible reconfiguration of RAN components, fostering an open and competitive ecosystem. This openness facilitates seamless integration of multi-vendor equipment, promotes innovation in network deployment, and supports the delivery of diverse and customizable user services.
@@
  • Intelligence: The ORAN architecture incorporates two RICs to enable intelligent network optimization. AI /ML algorithms can be efficiently employed to determine and implement policies and control actions at different timescales.
2.2 NTN
In this subsection, we review the basic background on NTN, including its components, payload types, and characteristics. A broad range of survey and magazine articles has examined NTNs from different perspectives. Several works focus on 3GPP standardization and evolution from 5G to 6G, including surveys [el2023introduction, azari2022evolution, rinaldi2020non], as well as magazine articles [hosseinian2021review, lin20215g, saad2024non, wigard2023ubiquitous]. Other works emphasize applications and system architectures, including surveys [openRAN-NTN-1, he2024non] and magazine papers [jamshed2025non]. In addition, several recent works, including [deng2025distributed, kaleem2024emerging, mahboob2024revolutionizing], investigate emerging technologies and intelligent NTN paradigms, with particular attention to AI-based solutions such as generative AI and reinforcement learning. These works complement the scope of this section and provide readers with a comprehensive overview of NTN research, covering technological evolution, standardization progress, and prospective future directions.
@@ 2.2.1 COMPONENTS
  • Geostationary Earth Orbit (GEO) satellites: GEO satellites typically maintain a fixed position at an orbit altitude of 35,786 km above the Earth’s equator. They provide extremely wide coverage and maintain fixed visibility from the ground, but they suffer from high propagation delay and significant path loss. GEO systems are commonly used for broadcasting, satellite television, and large-scale communication services that are less sensitive to latency [mahboob2025transforming].
@@
  • Medium Earth Orbit (MEO) satellites: MEO satellites are located at an orbit altitude of 2,000 to 25,000 km above the Earth. They offer a balance between coverage area, latency, and constellation size, achieving lower delays compared with GEO while requiring fewer satellites than Low Earth Orbit (LEO) satellite constellations to achieve global coverage [al2022survey]. MEO satellites are widely used in navigation and positioning systems such as Global Positioning System (GPS), China’s BeiDou navigation satellite system (BDS), Russia’s Global Navigation Satellite System (GLONASS), and the European Union’s Galileo [jiao2025geostationary].
@@
  • Low Earth Orbit (LEO) satellites: LEO satellites operate at altitudes of approximately 500 to 2,000 km above the Earth. This orbit altitude allows them to achieve a much shorter propagation delay and reduced path loss compared with a satellite system with a higher altitude. However, LEO satellites face several challenges, including their relatively smaller coverage, the need for frequent handovers, and the strong Doppler effect because of the high mobility [deng2025orthogonality]. In practice, typical LEO constellations such as Starlink have demonstrated real-world use in delivering global mobile broadband internet and improving connectivity in remote and underserved regions [hui2025review].
@@
  • High Altitude Platforms (HAPs): The HAPs, such as airplanes, balloons, and airships, usually reach the stratosphere region with an altitude ranging from 20 km to 50 km [shang2024enhancing, openRAN-NTN-1]. They combine the advantages of satellites and terrestrial systems by maintaining a quasi-stationary state, offering relatively wide coverage, low latency, and flexible deployment, which makes them typically used for civilian applications, such as providing broadband internet access to rural areas [liu2023joint].
@@
  • Uncrewed Aerial Vehicles (UAVs): UAVs typically operate at low altitudes ranging from 0.1 km to 15 km, and offer high mobility and deployment flexibility, which makes them suitable for rapidly providing temporary or emergency coverage in post-disaster scenarios [kaleem2024emerging]. However, UAVs suffer from limited endurance, rapidly varying communication channels, and complex interference management. Beyond emergency response, UAVs are also expected to play a central role in the emerging low-altitude economy (LAE), enabling services such as aerial logistics, infrastructure monitoring, and precision agriculture [ahmed2025toward].
2.2.2 PAYLOADS OPTIONS
@@
It is noted that two main types of payload architecture have been defined for NTN by 3GPP: transparent and regenerative [rinaldi2020non, mahboob2025transforming]. The detailed schemes behind these two architectures are introduced as follows:
@@
Transparent payload: With transparent payload architecture, the NTN platform simply relays the signal from users to the terrestrial gNodeB (gNB), which consists of CU, DU, and RU [wigard2023ubiquitous, saad2024non].
@@
  • Only RU: the NTN platform is equipped solely with the RU, responsible for basic RF functions or partial physical layer processing, depending on the selected functional split and NTN’s capabilities. DU and CU are deployed on the ground, resulting in long propagation delays on the wireless fronthaul link to RU [shah2025urllc].
  • Partial gNB (RU+DU): the NTN platform can take on some of the processing tasks that are normally performed by the full gNB on the ground. Then the signal will be further processed by the gNB-CU on the ground for further processing. This setting can improve the performance of the network, especially in areas with high traffic demand [rihan2023ran].
  • Full gNB (RU+DU+CU): the NTN platform would have complete processing capabilities to regenerate the signal in either the feeder link or the service link, which allows for better performance, such as lower latency and higher data rates [mahboob2025transforming].
2.3 AI-NATIVE FOR COMMUNICATION
@@
2.3.1 CONCEPT
Typically, the AI-based communication can be divided into two stages, i.e., AI-Aided and AI-Native [baek2023ai]. Specifically, an AI-Aided method aims to optimize specific communication modules, which has been shown to be practical and feasible in numerous studies [deng2025distributed, AI-for-TN-NTN-1, kaleem2024emerging, khan2023ai]. An AI-Native method enables the system to have intrinsic trustworthy AI capabilities, where AI is naturally embedded across system design, deployment, operation, and maintenance. Apart from that, AI-Native method implementation leverages a data-driven and knowledge-based ecosystem that continuously consumes and produces information to enable new AI functionalities or replace static, rule-based mechanisms with adaptive, learning-based intelligence when needed [iovene2023defining].
Figure 4: AI-Native architecture for communication based on [iovene2023defining].
Several recent studies and reports have discussed the concept of AI-Native communications and its applications in ORAN and NTNs [3GPP_AI_ML_NG_RAN, baek2023ai, khan2023ai, iovene2023defining]. The 3GPP report [3GPP_AI_ML_NG_RAN] discussed the integration of AI/ML in NG-RAN and 5G-Advanced systems as a step toward 6G evolution. The authors in [baek2023ai] outlined key challenges and future directions for realizing AI-native communication systems. [khan2023ai] analyzed AI-driven RAN design for 6G, highlighting major research issues, challenges, and potential solutions. Additionally, the Ericsson white paper [iovene2023defining] defined AI-Native as a key enabler for intelligent telecom networks and presented an architectural framework and maturity model.
2.3.2 ARCHITECTURE
Ericsson presented a concise and generic AI-Native architecture in [iovene2023defining]. Building upon this concept, we extend and refine its essential aspects to develop a detailed version, as illustrated in Fig. 4.
As shown in the figure, the architecture is structured around several core layers, including transport, cloud, access and mobility, management and orchestration, and applications. These layers interact closely with distributed intelligence and data resources, enabling seamless execution and adaptation of AI models across different domains and deployment environments. It is noted that the cross-domain model lifecycle management helps enable coordinated and trusted intelligence, constantly improving and following data changes, to achieve system-wide end-to-end performance gains.
The AI-Native architecture is characterized by four fundamental aspects [iovene2023defining]:
@@ 2.3.2 ARCHITECTURE
  • Distributed Data Infrastructure: This aspect enables the data and knowledge sharing across different layers. It supports the aspect of intelligence everywhere by ensuring the availability of data and computing resources, and also interacts with the model orchestrator for the transportation of data or intelligence. For example, under stringent requirements on the timing of the data, it is often more efficient to move intelligence closer to the data source, rather than transporting data to intelligence.
@@
  • AI as a Service (AIaaS): Certain AI functionalities, such as AI model lifecycle management features and data handling, can be exposed as services to external parties. This concept, known as AIaaS, transforms networks into open innovation platforms where service providers and even end users can flexibly access and utilize AI resources on demand.
@@ 3 NTN DEVOPS CHALLENGES AND SOLUTIONS
Different from traditional TNs, the DevOps lifecycle in NTN faces multiple unique challenges, where a typical product DevOps lifecycle is given in Fig. 5 [herrera2025tutorial, Corici2023]. Therefore, in this section, we focus on analyzing unique challenges in the DevOps lifecycle of NTN and the corresponding solutions via integrating ORAN principles.
@@
3.1 NTN PLATFORM SWAP LIMITATION

### v3

@@ 2 BACKGROUND
This section is intended to provide a brief background about ORAN, NTNs, and AI-Native for communication. We have listed some important and representative survey and/or magazine papers in each subsection for the interested readers to further explore these topics.
2-A ORAN
In this sub-section, we aim to provide the readers with a clear understanding of the ORAN architecture and its key characteristics. A wide range of survey and tutorial articles have examined ORAN from various perspectives, providing valuable insights into its architecture, design principles, and implementation. Representative works [99, 100, 21, 128, 10] offered comprehensive overviews of ORAN development, interfaces, and open research challenges. In addition, other surveys and tutorials presented a broad overview of ORAN along with focused discussions on specific topics such as network slicing [15], simulation and testbeds [57], and spectrum allocation [81]. The aforementioned surveys and tutorials can serve as valuable complementary resources for readers seeking further background on ORAN relevant to this study.
2-A.1 ARCHITECTURE
The RAN architectures have evolved from the conventional designs used in 2G to 4G networks toward virtualized RAN (vRAN) and cloud RAN (C-RAN), in which network functions and services are decoupled from dedicated hardware and implemented virtually [25]. This evolution has further led to ORAN, where multi-vendor components are disaggregated, virtualized, and connected through open and well-defined interfaces.
The O-RAN Alliance was founded in 2018 to promote openness and intelligence in the RAN after the two organizations C-RAN Alliance and the xRAN forum joined forces. The main goal of the Alliance is to standardize an architecture and set of interfaces for the ORAN paradigm [127]. The alliance introduces the ORAN architecture and specifications to extend the RAN standards to include openness and intelligence. The alliance has 11 technical work groups (WGs) for the specification tasks in the ORAN architecture, 5 focus groups for the topics that are over-arching the technical WGs or are relevant for the whole organization, and a next Generation Research Group (nGRG) focusing on the research of open and intelligent RAN principles in 6G and future network standards [16].
@@ 2-A.1 ARCHITECTURE
  • SMO: SMO provides the end-to-end (E2E) lifecycle management, such as the orchestration and fault, configuration, accounting, performance, and security (FCAPS) operations, of the ORAN network functions and RICs [24].
  • O-eNB: O-eNB may take the form of either a legacy eNB or an ng-eNB, since the ORAN architecture also supports the integration of LTE base-stations.
@@
  • Y1 Consumer: Y1 Consumer refers to the entity within or outside of the public land mobile network (PLMN) trust domain that consumes the Y1 services produced by the near-real-time RIC (Near-RT RIC) [89].
ORAN architecture defines disaggregated functional entities, including the ORAN radio unit (O-RU), ORAN distributed unit (O-DU), and ORAN centralized unit (O-CU), which are interconnected through standardized open interfaces specified by the O-RAN Alliance and 3GPP. The O-RAN Alliance and 3GPP has specified different split options as discussed later in Section 4. Following the typical functional split Option 7.2x [70, 91], the main functionalities of these entities are:
@@
In addition, two types of RICs, including non-real-time RIC (Non-RT RIC) and Near-RT RIC, enable the intelligent RAN optimization across different layers and timescales [92, 91]. Their key functionalities are summarized as follows:
@@
To provide essential background on the interfaces defined by 3GPP but adopted by the O-RAN Alliance [91, 15], brief descriptions are presented as follows:
@@
Figure 2: The latest ORAN Architecture by O-RAN Alliance [91].
Furthermore, we introduce the new interfaces defined by the O-RAN Alliance below [99, 91].
@@
  • Y1: Y1 interface enables the Y1 consumers to subscribe or request the RAN analytics information provided by Near-RT RIC [89].
  • D2: D2 interface is newly defined as the interface between two O-DU connected to the same O-CU-CP, and supports NR carrier aggregation of carriers configured in different O-DU nodes [90].
2-A.2 KEY FEATURES
@@
  • Open Interfaces: The open interface design enables full interoperability and flexible reconfiguration of RAN components, supporting an open and competitive ecosystem. This openness facilitates seamless integration of multi-vendor equipment, promotes innovation in network deployment, and supports the delivery of diverse and customizable user services.
@@
  • Intelligence: The ORAN architecture incorporates two RICs to enable intelligent network optimization. AI /ML algorithms will be more efficiently employed to determine and implement policies and control actions at different timescales.
2-B NTN
In this subsection, we review the basic background on NTN, including its components, payload types, and characteristics. A broad range of survey and magazine articles has examined NTNs from different perspectives. Several works focus on 3GPP standardization and evolution from 5G to 6G, including surveys [43, 20, 104], as well as magazine articles [59, 74, 105, 129]. Other works emphasize applications and system architectures, including surveys [87, 55] and magazine papers [64]. In addition, several recent works, including [38, 67, 79], investigate emerging technologies and intelligent NTN paradigms, with particular attention to AI-based solutions such as generative AI and reinforcement learning. These works complement the scope of this section and provide readers with a comprehensive overview of NTN research, which covers both technological evolution and standardization progress.
@@ 2-B.1 COMPONENTS
  • Geostationary Earth Orbit (GEO) satellites: GEO satellites typically maintain a fixed position at an orbit altitude of 35,786 km above the Earth’s equator. They provide extremely wide coverage and maintain fixed visibility from the ground, but they suffer from high propagation delay and significant path loss. GEO systems are commonly used for broadcasting, satellite television, and large-scale communication services that are less sensitive to latency [78].
@@
  • Medium Earth Orbit (MEO) satellites: MEO satellites are located at an orbit altitude of 2,000 to 25,000 km above the Earth. They offer a balance between coverage area, latency, and constellation size, achieving lower delays compared with GEO while requiring fewer satellites than Low Earth Orbit (LEO) satellite constellations to achieve global coverage [13]. MEO satellites are widely used in navigation and positioning systems such as Global Positioning System (GPS), China’s BeiDou navigation satellite system (BDS), Russia’s Global Navigation Satellite System (GLONASS), and the European Union’s Galileo [65].
@@
  • Low Earth Orbit (LEO) satellites: LEO satellites operate at altitudes of approximately 500 to 2,000 km above the Earth. This orbit altitude allows them to achieve a much shorter propagation delay and reduced path loss compared with a satellite system with a higher altitude. However, LEO satellites face several challenges, including their relatively smaller coverage, the need for frequent handovers, and the strong Doppler effect because of the high mobility [37]. In practice, typical LEO constellations such as Starlink have demonstrated real-world use in delivering global mobile broadband internet and improving connectivity in remote regions [61].
@@
  • High Altitude Platforms (HAPs): The HAPs, such as airplanes, balloons, and airships, usually reach the stratosphere region with an altitude ranging from 20 km to 50 km [111, 87]. They combine the advantages of satellites and terrestrial systems by maintaining a quasi-stationary state. Meanwhile, they offer relatively wide coverage, low latency, and flexible deployment. These characteristics make them well suited for civilian applications, such as providing broadband internet access to rural areas [75].
@@
  • Uncrewed Aerial Vehicles (UAVs): UAVs typically operate at low altitudes ranging from 0.1 km to 15 km. Due to their high mobility and deployment flexibility, they are suitable for rapidly providing emergency coverage in post-disaster scenarios [67]. However, UAVs suffer from limited endurance, rapidly varying communication channels, and complex interference management. Beyond emergency response, UAVs are also expected to play a central role in the emerging low-altitude economy (LAE), enabling services such as aerial logistics, infrastructure monitoring, and precision agriculture [11].
2-B.2 PAYLOAD OPTIONS
@@
There are typically two main types of payload architecture defined for NTN by 3GPP: transparent and regenerative [104, 78]. The detailed schemes behind these two architectures are introduced as follows:
@@
Transparent payload: With transparent payload architecture, the NTN platform simply relays the signal from users to the terrestrial gNodeB (gNB), which consists of CU, DU, and RU [129, 105].
@@
  • Only RU: the NTN platform is equipped solely with the RU, responsible for basic RF functions or partial physical layer processing, depending on the selected functional split and NTN’s capabilities. DU and CU are deployed on the ground, resulting in long propagation delays on the wireless fronthaul link to RU [110].
  • Partial gNB (RU+DU): the NTN platform can take on some of the processing tasks that are normally performed by the full gNB on the ground. Then the signal will be further processed by the gNB-CU on the ground for further processing. This setting can improve the performance of the network, especially in areas with high traffic demand [103].
  • Full gNB (RU+DU+CU): the NTN platform would have complete processing capabilities to regenerate the signal in either the feeder link or the service link, which allows for better performance, such as lower latency and higher data rates [78].
2-C AI-NATIVE FOR COMMUNICATION
@@
Figure 4: AI-Native architecture based on [63].
2-C.1 CONCEPT
Typically, the AI-based communication can be divided into two stages, i.e., AI-Aided and AI-Native [22]. Specifically, an AI-Aided method aims to optimize specific communication modules, which has been shown to be practical and feasible in numerous studies [38, 28, 67, 69]. An AI-Native method will enable the system to have intrinsic trustworthy AI capabilities, where AI is naturally embedded across system design, deployment, operation, and maintenance. In addition, AI-native implementations leverage a data-driven and knowledge-based ecosystem that continuously consumes and produces information. This ecosystem can help enable new AI functionalities, or replace the static and rule-based mechanisms with adaptive and learning-based intelligence when needed [63].
Several recent studies and reports have discussed the concept of AI-Native communications and its applications in ORAN and NTNs [3, 22, 69, 63]. The 3GPP report [3] discussed the integration of AI/ML in NG-RAN and 5G-Advanced systems as a step toward 6G evolution. The authors in [22] outlined key challenges and future directions for realizing AI-Native communication systems. In [69], the authors analyzed AI-driven RAN design for 6G, highlighting major research issues, challenges, and potential solutions. Additionally, the Ericsson white paper [63] defined AI-Native as a key enabler for intelligent telecommunication networks and presented an architectural framework and maturity model.
2-C.2 ARCHITECTURE
Ericsson has presented a concise and generic AI-Native architecture in [63]. Building upon this concept, we extend and refine its essential aspects to develop a detailed version, as illustrated in Fig. 4.
As shown in the figure, the architecture is structured around several core layers, including transport, cloud infrastructure, access and mobility, management and orchestration, and applications. These layers interact closely with distributed intelligence and data resources to support seamless execution and adaptation of AI models across different domains. Moreover, the cross-domain model lifecycle management helps enable coordinated and trusted intelligence, constantly improving and following data changes, to achieve system-wide E2E performance gains.
The AI-Native architecture is characterized by four fundamental aspects [63]:
@@ 2-C.2 ARCHITECTURE
  • Distributed Data Infrastructure: This aspect enables the data and knowledge sharing across different layers. It supports the aspect of intelligence everywhere by ensuring the availability of data and computing resources. In addition, it also interacts with the model orchestrator for the transportation of data or intelligence. For example, under stringent requirements on the timing of the data, it is often more efficient to move intelligence closer to the data source, rather than transporting data to intelligence.
@@
  • AI as a Service (AIaaS): Certain functionalities, such as AI model training or execution environments, and data handling, can be exposed as services to external parties. This concept, known as AIaaS, transforms networks into open innovation platforms, where service providers and even end users can flexibly access and utilize AI resources on demand.
@@ 3 NTN DEVOPS CHALLENGES AND SOLUTIONS
Different from traditional TNs, the DevOps lifecycle in NTN faces multiple unique challenges [57, 33], as illustrated in Fig. 5. In this section, we focus on analyzing these DevOps challenges of NTN and the corresponding solutions via integrating ORAN principles.
@@
Table 3: Typical specifications of NTN platforms [38].
Specification Untethered UAV Tethered UAV HAP LEO MEO GEO
Loading Capability 2.7 kg 15 kg 140 kg - - -
Power Supply 0.3 kWh 30 kWh 20 kWh - - -
Latency 1 ms 1 ms 1 ms 30-50 ms 150 ms 600 ms
Endurance 30 mins 24 hours 2 months 5-7 years 10-15 years 15-20 years
Deployment Cost Low Low Medium High High High
Network Type 1 Micro/Pico 1 Macro 7 Multiple Macro 61 Multiple Macro 61 Multiple Macro 61 Multiple Macro
3-A NTN SPECIFICATIONS AND CONSTRAINTS